OpenStudy (anonymous):

solve the equation tan(theta) = 1 where 0

OpenStudy (anonymous):

do u want to know sin,cos,cot....etc?

OpenStudy (anonymous):

if u want to then u have the identity tan^2(theta) +1=sec^2(theta) this is the first thing u need .the second thing is that u should know is at which quarter the theta will be this can be known from the question bcuz tan(theta) =1 that mean it a positive quantity which mean it either in first quarter or third quarter u suppose now the theta first in first quarter & apply the identity to get sec and u can also know the cos from sec then u apply the other identity that say sin^2=1-cos^2 after u find the other triangular function u suppose the theta in the third quarter and do the same thing as in first one